OpenStudy (anonymous):

Which functions are correct? f(x)=2^4x g(x)=2^x Choose exactly two answers that are correct. A. (f g)(x)=2^5x B. (f g)(x)=2^4x2 C. (f/g)(x)=2^3x D. (f/g)(x)=1/2^4x

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

Wait nevermind It worked!

OpenStudy (anonymous):

So what is the next step to do??

jigglypuff314 (jigglypuff314):

so now that you know how those equations work we know that (f * g)(x) = 2^(4x) * 2^(x) right?

OpenStudy (anonymous):

Yes

jigglypuff314 (jigglypuff314):

can you simplify that? which answer do you think that would be?

OpenStudy (anonymous):

I think it would be B

jigglypuff314 (jigglypuff314):

hmmm do you know about these properties? https://mathtestpreparation.com/Principles/algebra_sproduct2.gif

OpenStudy (anonymous):

Yes!

jigglypuff314 (jigglypuff314):

so 2^(4x) * 2^(x) would equal 2^(4x + x) = ?

OpenStudy (anonymous):

no wait is it 16, I am so confused

jigglypuff314 (jigglypuff314):

4x + x = ?

OpenStudy (anonymous):

4x

jigglypuff314 (jigglypuff314):

are you sure? ^_^

OpenStudy (anonymous):

no 5x

OpenStudy (anonymous):

I think I am sure

jigglypuff314 (jigglypuff314):

good so 2^(4x) * 2^(x) = 2^(4x + x) = 2^(5x) do you see how that works?

OpenStudy (anonymous):

Yes I do

jigglypuff314 (jigglypuff314):

now, what do you think the answer would be then?

OpenStudy (anonymous):

the answer would be A, but it says we have to choose to answers

jigglypuff314 (jigglypuff314):

indeed the other answer would be the idea of (f / g)(x) = 2^(4x) / 2^(x)

OpenStudy (anonymous):

So would it be A and B

jigglypuff314 (jigglypuff314):

nono, we just got A right? now we are looking at the dividing idea (between C and D)

jigglypuff314 (jigglypuff314):

(f / g)(x) = 2^(4x) / 2^(x) = 2^(4x - x) = ?

OpenStudy (anonymous):

hmm 15...

OpenStudy (anonymous):

Wait, the answers are A and C

jigglypuff314 (jigglypuff314):

there you go

OpenStudy (anonymous):

Thank you so much!!!!!!!!